solve the systme by using an augmented matrix
4x+y-z=7
-2x+2y+5z=3
7x-3y-9z=-4

Answer by Alan3354(21608)

4x+y-z=7
-2x+2y+5z=3
7x-3y-9z=-4
-------------
x = -1
y = 8
z = -3
I have an Excel sheet that does these. If you have Excel and want a copy, email me.
Doing them by hand is tedious and error prone. It's important to know how to do them, but once you know, it's better to avoid it.